answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many ways you can arrange 6 cars in equal rows and columns?

If you include a row or column of 1, the answer is 4: 1 by 6, 2 by 3, 3 by 2, and 6 by 1.

What is probability of 1 even die and 1 odd when rolling 2 dice?

The probability is 1/2.

The probability is 1/2.

The probability is 1/2.

The probability is 1/2.

What is the probability of independent events?

If you mean the probability that independent events will occur, it is the product of the individual probabilities that they occur. Ex: Probability of picking a red marble from a jar with 4 red and 6 green marbles and picking a black card from a deck of 52.

P(red) = 4/10 = 2/5

P(black card0 = 26/52 = 1/2

P(both) = 2/5 x 1/2 = 1/5

If you roll two dice what is the probability of the event that you roll a total of seven or ten on the two dice?

To roll a 7: 1/6, 2/5, 3/4, 4/3, 5/2, 6/1 To roll a 10: 4/6, 5/5, 6/4 total of 9 chances out of 36 or 1 in 4

If the football team that wins the coin toss elects to defend a goal who gets the ball in the second half?

The team that lost the coin toss at the beginning of the game would get first choice to start the second half. It would be their choice to defend a goal, kickoff, or receive the ball.

What is the size of the sample space for all the outcomes possible from rolling four dice?

There is 6 possible outcomes per roll of a die. So, there are 6*6*6*6 outcomes or 64 or 1296 possible outcomes.

What are the odds of being an albino?

The worldwide probability of albinism is about 1 in 17,000, or about 0.0000588. Prevalence varies, and is highest among sub-Saharan African people.

For more information, please see the Related Link below.

What is the probability of selecting a constant from the letters?

Until the letter is selected, it is a variable. Immediately after it is selected, the outcome is no longer a variable but a constant.

What happens if a value falls below the mean 2 standard deviation when the probability to belong to the distribution is lower than 0.05?

Nothing actually happens! You just get a value that is very unlikely but still possible. That it is possible is evidenced by the fact that the value was observed.

What are the limitation of the probability?

Probability of an even must lie in the closed interval [0, 1].

Probability of an even must lie in the closed interval [0, 1].

Probability of an even must lie in the closed interval [0, 1].

Probability of an even must lie in the closed interval [0, 1].

What is the chance of getting an even number when you toss a number cube 10 times?

It is

1 - Prob(No even numbers in 10 throws)

= 1 - [Prob(Not an even number)]10 = 1 - (1/2)10 = 1 - 1/1024 = 1023/1024 or 0.9990

So you are almost certain to get an even number.

1928 Indian head penny?

Check again. The last year for Indian-head pennies was 1909.

The amount of cards in a standard deck?

Fifty-two (52), not including two jokers included in most decks.

What happens in the event of a vacancy occurring in either chamber of congress?

Whenever a seat in the U.S. Senate or House of Representatives is vacated due to death, resignation or removal, the State Governor issues a Writ of Election, and a special election is held to fill the seat for the remainder of the term. The special election can coincide with a regular congressional election. In the case of a vacated Senate seat, the Governor can appoint someone to temporarily fill the seat until an election can be held if the state legislature grants him that power.

If a card is drawn from a deck numbered 1 through 10 and a die is rolled what is the probability of drawing a 10 and a 3?

The probability of drawing the 10 is 1/10 and the probability of rolling a 3 is 1/6. So, the probability of both is 1/10 * 1/6 = 1/60.

What is the probability of getting a 7 in a pack of cards?

It is 1: a certainty. A complete pack of cards must have a seven.

Trending Questions
What is the probability of selecting a red picture card from a deck of cards? If rolling three dice what is the probability of rolling three ones? What is the the relationship between population and sample parameter and statistic? Will learning about Probability Binomial and Poisson Probability Distributions using Excel make our studies easier? What are some limitations of statistics? What is percentile deviation? How many ways can you roll number 6 with 2 dice? What is the probability a student will get 8 out of 10 right on a true or false test if he quesses on each question? What is the meaning of expected outcomes? I have a fair air six sided dice. numbered one to six. What is the probability that you roll a number less than five? What is the theoretical probability of spinning a 3 and picking an a? How can you use theoretical probability to make predictions in real-world situations? What are the adjectives of probability? If you select two cards from a deck of 52 well shuffled playing cards what is the probability that both cards will be black? In probability a trial where you dont know the outcome? What would be the experimental probability of rolling a die 50 times and getting 4? What is Triamterene-h? Each different letter of the word monomial is written on a card One card is drawn How many possible outcomes are there? What is the probability of not obtaining either an ace or a king in a single blind draw from standard deck of card? The probability of the contractor failing to meet the requirements of the potential contract and rdquo is considered to be what type of risk?